In a large plant that has life of 12 years, the management is considering two different pumps. Which of pumps should they purchase? The data on the pumps is given in the following Table. Minimum attractive interest rate is 15%. You can use net present worth (cost) criteria with common multiple life which is 6. Do before tax analysis which means do not account for tax and depreciation. Pump TCI (S) Operating Cost Life of alternatives 100,000 150,000 (S/year) 50,000 40,000 years) 2
Answer #1

Alternative A

Initial cost = $100,000

Operating cost = $50,000 per year

Life = 2 years

The common multiple life is 6 years. This means this alternative will be replaced by similar pump for two times after year 2 and year 4.

Calculate the net present worth -

NPW = -100,000 - 50,000 (P/A, 15%, 6) - 100,000 (P/F, 15%, 2) - 100,000 (P/F, 15%, 4)

NPW = -100,000 - (50,000 * 3.7845) - (100,000 * 0.7561) - (100,000 * 0.5718)

NPW = -100,000 - 189,225 - 75,610 - 57,180

NPW = -422,015

The net present worth of Alternative A is -$422,015

Alternative B

Initial cost = $150,000

Operating cost = $40,000 per year

Life = 3 years

The common multiple life is 6 years. This means this alternative will be replaced by similar pump for one time after year 3.

Calculate the net present worth -

NPW = -150,000 - 40,000 (P/A, 15%, 6) - 150,000 (P/F, 15%, 3)

NPW = -150,000 - (40,000 * 3.7845) - (150,000 * 0.6575)

NPW = -150,000 - 151,380 - 98,625

NPW = -400,005

The net present worth of Alternative B is -$400,005

The net present worth of Alternative B is numerically higher.

So,

The company will purchase Alternative B.